# The SST-1M imaging atmospheric Cherenkov telescope for gamma-ray astrophysics11footnotetext: Web page: https://sst-1m.space

### v1

@@ 2.1 Optical designs for IACTs
IACT optical designs are typically based on the Davies-Cotton or parabolic designs. Originally, the D-C design was invented for the optics of a solar concentrator [4] as this system reduces the size of off-axis optical aberrations for field of views (FoVs) of several degrees. These aberrations have a lower value in comparison to a parabolic mirror with identical aperture, defined as $f/D$, where $D$ is the optical surface diameter and $f$ the focal length. The reason why the D-C is adopted in IACTs is related to cost and ease of maintenance in the field, as they function without a protective dome. The reflector of a D-C telescope is tessellated into identical elements with the same radius of curvature $R$, shape and size arranged on a spherical surface (dish) with a radius of R///2. The orientation of the mirror segments is such that the light from the point-like source at infinity is reflected to the focal plane of the D-C setup. The focal plane of the telescope is located at $f^{\prime}=R/2$ (see Fig. 1). Identical segments with a spherical reflective surface are easy to manufacture and can be easily adjusted or replaced when necessary. In contrast, solutions, such as parabolic mirrors, cannot use identical segments for the entire surface.
@@ 2.2 The SST-1M main features
The physics objectives of the SSTs span the energy range from $500$ GeV to beyond $100$ TeV. The sensitivity at the highest energies, which is critical for exploring the Galaxy, depends on the extension of the light pool on the ground and on the number of synchronised telescopes forming the array to sample it and on their mutual distance. These physics goals include large-scale surveys, high-precision spectral measurements in the high-energy range to understand cosmic ray acceleration to PeV energies (the energy of the ‘knee’ in the cosmic ray spectrum), and the search for cosmic ray sources. It is then fundamental to achieve the longest possible exposure time to collect the poor gamma-ray statistics above 10 TeV and also to achieve designs with wide FoV, for both facilitating aperture overlaps between telescopes and observations of the many interesting high-energy extended galactic sources.
@@
The SiPM sensors are lightweight, consume roughly an order of magnitude less power than photomultiplier tubes (PMTs), and offer high photodetection efficiency (PDE). However, they have the drawback of higher noise, which is influenced by their signal duration and capacitance, which is proportional to size. SiPM provide a PDE that is 10-20 % higher than that of comparable PMTs. Aditionally, their sensitivity surpasses that of the typical PMTs beyond 600 nm and extends to wavelengths above $800$ nm, unlike PMTs. While their higher dark count rate than PMTs at room temperature is not a major concern as NSB dominates, their higher sensitivity in the red and near-infrared regions requires the adoption of filters on the camera entrance window or the lightguides / sensors. As the NSB increases in this wavelength region, the SST-1M Borofloat entrance window is coated with a dichroic filter cutting wavelengths longer than 550 nm. Interestingly, the SST-1M has a DC-coupled preamplifier camera electronics, unlike FACT which is AC-coupled. This configuration allows for the monitoring of the NSB through changes of the baseline.
@@
Figure 5: The camera of an SST-1M at the Ondřejov site.
@@
Figure 6: Table of main SST-1M parameters. *) The mirror effective area takes the reflectivity of mirrors and the telescope structure shadowing into account.
Figure 7: The first (left) and second (right) SST-1M installed at Ondřejov.
@@ 3.1.2 The drive system
Figure 14: The IMO slew drive. Reproduced with permission of IMO Holding GmbH.
@@ 3.4.1 The photo-detection plane
Figure 24: Blow-up CAD of a PDP module (left) and its picture partially assembled (right). There are 108 modules in the PDP each weighting 300 g.
@@
To have a spatial uniform response of the camera, the pixels should have a circular shape to ensure the same distance between their centres in every direction. Nonetheless, a circular shape would introduce large dead spaces, and so the selected pixels are hexagonal. A hexagonal SiPM device, named S10943-2832(X), has been designed in collaboration with Hamamatsu and is shown in Fig. 25. This device is based on the so-called low cross-talk technology LCT2, which was available when the camera design was done. Currently, the next generation of devices is on market such as LCT5 (S13360) and LVR (S13360). The LCT5 offers better performance than LCT2 in terms of PDE and cross talk. The sensor area is around 93.6 mm2 with linear size of 9.3 mm. The Winston cone design, shown on the right of Fig. 25, is also described in [11]. The sensor capacitance is directly related to its active area and this has an impact on the signal decay time and consequently on the read-out electronics noise.
@@
The adopted solution is a trans-impedance amplifier for its high bandwidth. Given the small size of the 12-pixel-module PCB and other constraints in terms of power and cost, it is not possible to have an amplifier for each of the 4 channels of the sensors. The solution is summing two channel signals at the input of one low noise amplifier, so that two low noise amplifiers are needed per pixel. The sum of the output signals is made in a differential stage which guarantees a good single photo-electron resolution with the LCT2.
@@ 4.2.2 The optical point spread function of the telescope
The PSF represents the optical quality of the telescope. We define the PSF as the angular diameter of a circle containing 80% of the photons emitted by a point source, denoted by D80. The centre of the circle is chosen to be the centre of gravity of the photon distribution. Given that the spot profile at the focal plane is not necessarily symmetrical, we define D80 by starting from the centre of gravity of the photon distribution and integrating the signal in larger and larger circles until 80% of the total emission is reached. For the camera field of view of 9, the PSF needs to be determined for light rays coming from a source located at $\sqrt{0.8}\times 4.5^{\circ}$, i.e. at 4 off-axis. The PSF simulation for the various field of view of the telescope using two independent simulation tools is shown in Fig. 4. The simulation is compared with the real measured PSF of the telescope in Sec. 2 and it is shown in 2 dimensions in Fig. 30.
@@ 4.3.1 Prototype structure tests
To fully identify the spectrum of natural modes of the telescope structure, additional modal experiments were conducted using piezoelectric accelerometers by a specialised team from AGH in Krakow. Two accelerometers were used – one at the bottom of the dish support structure and another on the back of the camera box. The structure was excited with a PCB 086D20 impact hammer at various points on the frame and different directions. Points of excitation were approximately the same across different elevation angles ($-10^{\circ}$ – close to parking position, but telescope not locked in docking station, $-13^{\circ}$ – telescope locked, $45^{\circ}$ and $90^{\circ}$), except near the camera box at $45^{\circ}$ and $90^{\circ}$ angles. Nineteen modal experiments were performed, varying the excitation points (dish support, mast, camera box, counterweight) and directions ($x$, $y$, $z$). Some measurements were conducted with the drives activated (typical during observation) and others with the drives turned off and brakes engaged (as in the parking position). The reference frame was centered at the base of the telescope tower, with the $x$-axis aligned along the telescope mast and the $y$- and $z$-axes oriented vertically. The analysis was based on spectral transfer functions and coherence functions. The spectral transfer function provides information about the natural frequencies excited during an experiment, while the coherence function indicates the correlation between the excitation and the response, ranging from 0 to 1. Results are considered valid when the coherence function values exceed 0.8. Frequencies analyzed ranged from 1.9 Hz to 32 Hz.
@@ 4.4.2 Telescope optical efficiency
We calculate the fraction of the Cherenkov light signal that can be detected by the telescopes using the calculated total efficiency from 290 nm to 700 nm: 15.1% and 16.5% for the prototype and the second telescope. The same efficiencies for the NSB are 12.2% and 13.6%. The resulting SNR ratio is shown in Fig. 35 (bottom). One can clearly see the impact of the low-pass optical filter applied on the entrance window, which stabilises the SNR ratio above 540 nm.
Figure 35: Efficiency of the optical telescope elements dedicated to transmission of light from the mirror dish (left). The total efficiencies of the 2 telescopes are 15.1% and 16.5%. SNR for the 2 telescopes (right).
@@ 4.4.3 Sensor characteristics
Gain [ADC] DCR [MHz] OCT [%] $\sigma_{pe}$ [ADC] $\sigma_{el}$ [ADC]
@@
Table 9: Summary of mean and standard deviation of the distributions of the main SiPM parameters. DCR refers to dark count rate, OCT, to optical cross talk.
@@ 4.6 Operation at the Ondřejov site
Figure 43: The location of the two SST-1Ms at the Ondřejov observatory (left). The layout of the telescope surroundings. The telescope is secured with a fence accessible via a gate. The camera servers and all necessary control electronics are located in a wooden booth at the top right of the image on the right image. The DAQ and slow control are located in the control cabinet next to the chiller.
@@ 5 Summary
The high-performance, large-area SiPMs with dedicated slow control, their lightweight (8.6 t) and compact mechanical telescope structure and its robotic operation require minimal maintenance in harsh environment, and make the SST-1M suitable for for deployment at high-altitude sites above 4000 m a.s.l. The SiPM market and their applications have widely expanded in the automotive, biology, medical imaging and LIDAR fields. Major companies such as Broadcom, Hamamatsu, Ketek, On-Semi, Advansid, and FBK are highly competitive and rapidly reducing costs.
@@ Acknowledgments
This publication was created as part of the projects funded in Poland by the Minister of Science based on agreements number 2024/WK/03 and DIR/WK/2017/12. The construction, calibration, software control and support for operation of the SST-1M cameras is supported by SNF (grants CRSII2_141877, 20FL21_154221, CRSII2_160830, _166913), by the Boninchi Foundation and by the Université de Genève, Faculté de Sciences, Départment de Physique Nucléaire. The Czech partner institutions acknowledge support of the infrastructure and research projects by Ministry of Education, Youth and Sports of the Czech Republic and regional funds of the European Union, MEYS LM2023047 and EU/MEYS CZ.02.01.01/00/22 008/0004632, and Czech Science Foundation, GACR 23-05827S.
Appendix A Appendix section
A.1 Telescope and infrastructure diagrams

### v2

@@ 2.1 Optical designs for IACTs
IACT optical designs are typically based on the Davies-Cotton or parabolic designs. Originally, the D-C design was invented for the optics of a solar concentrator [4] as this system reduces the size of off-axis optical aberrations for field of views (FoVs) of several degrees. These aberrations have a lower value in comparison to a parabolic mirror with identical aperture, defined as $f/D$, where $D$ is the optical surface diameter and $f$ the focal length. The reason why the D-C is adopted in IACTs is related to cost and ease of maintenance in the field, as they function without a protective dome. The reflector of a D-C telescope is tessellated into identical elements with the same radius of curvature $R$, shape and size arranged on a spherical surface (dish) with a radius of $\frac{R}{2}$, which also defines the focal distance from the detection plane, namely the camera (see Fig. 1). The orientation of the mirror segments is such that the light from the point-like source at infinity is reflected to the focal plane of the D-C setup. Identical segments with a spherical reflective surface are easy to manufacture and can be easily adjusted or replaced when necessary. In contrast, solutions, such as parabolic mirrors, cannot use identical segments for the entire surface.
@@ 2.2 The SST-1M main features
The physics objectives of the SSTs span the energy range from $500$ GeV to beyond $100$ TeV. The sensitivity at the highest energies, which is critical for exploring the Galaxy, depends on the extension of the light pool on the ground and on the number of synchronised telescopes forming the array to sample it and on their mutual distance. These physics goals include large-scale surveys, high-precision spectral measurements in the high-energy range to understand cosmic ray acceleration to PeV energies (the energy of the ‘knee’ in the cosmic ray spectrum), and the search for cosmic ray sources. To effectively collect the sparse gamma-ray data above 10 TeV, it is essential to deploy arrays of numerous telescopes spanning the largest possible area. Each telescope should also maximize its exposure time and field of view (FoV). This approach enables better aperture overlaps between telescopes and facilitates observations of the many intriguing high-energy, extended galactic sources.
@@
The SiPM sensors are lightweight, consume roughly an order of magnitude less power than photomultiplier tubes (PMTs), and offer high photodetection efficiency (PDE). However, they have the drawback of higher noise, which is influenced by their signal duration and capacitance, which is proportional to size. SiPM provide a PDE that is 10-20 % higher than that of comparable PMTs. Additionally, their sensitivity surpasses that of the typical PMTs beyond 600 nm and extends to wavelengths above $800$ nm, unlike PMTs. While their higher dark count rate than PMTs at room temperature is not a major concern as NSB dominates, their higher sensitivity in the red and near-infrared regions requires the adoption of filters on the camera entrance window, or lightguides, or sensors. As the NSB increases in this wavelength region, the SST-1M Borofloat entrance window is coated with a dichroic filter cutting wavelengths longer than 550 nm (see Sec. 3.4.1 and [14]). Interestingly, the SST-1M has a DC-coupled preamplifier camera electronics, unlike FACT which is AC-coupled. This configuration allows for the monitoring of the NSB through changes of the baseline.
@@
Figure 5: The camera of an SST-1M at the Ondřejov site. The reflectance in the yellow-red of the narrow-band filter on the camera window is visible.
@@
*) The mirror effective area takes the reflectivity of mirrors and the telescope structure shadowing into account.
Figure 6: Table of main SST-1M parameters.
Figure 7: The first (top) and second (bottom) SST-1M installed at Ondřejov.
@@ 3.1.2 The drive system
Figure 14: The IMO slew drive model WD-H. Reproduced with permission of IMO Holding GmbH. This figure is not publicly available online and should be requested on the manufacturer website: https://www.imo.de/en/products/slew-drives.
@@ 3.4.1 The photo-detection plane
Figure 24: Blow-up CAD of a PDP module (top) and its picture partially assembled (bottom). There are 108 modules in the PDP each weighting 300 g.
@@
To have a spatial uniform response of the camera, the pixels should have a circular shape to ensure the same distance between their centres in every direction. Nonetheless, a circular shape would introduce large dead spaces, and so the selected pixels are hexagonal. A hexagonal SiPM device, named S10943-2832(X), has been designed in collaboration with Hamamatsu and is shown in Fig. 25. This device is based on the so-called low cross-talk technology LCT2, which was available when the camera design was done. Currently, the next generation of devices is on market such as LCT5 (S13360) and LVR (S13360). The LCT5 offers better performance than LCT2 in terms of PDE and cross talk. The sensor area is around 93.6 mm2 with linear size of 9.3 mm. The lightguide design, shown on the right of Fig. 25, is also described in [11], where their reflectance is also reported on. The sensor capacitance is directly related to its active area and this has an impact on the signal decay time and consequently on the read-out electronics noise.
@@
The adopted solution is a trans-impedance amplifier for its high bandwidth. Given the small size of the 12-pixel-module PCB and other constraints in terms of power and cost, it is not possible to have an amplifier for each of the 4 channels of the sensors. The solution is summing two channel signals at the input of one low noise amplifier, so that two low noise amplifiers are needed per pixel. The sum of the output signals is made in a differential stage which guarantees a good single photo-electron resolution with the LCT2. In order to achieve the required dynamic range, the pre-amplifying stage saturates at around 600 p.e. Beyond this number of detected photons, the response is not linear, the amplitude saturates while the pulse integral continues to grow. A transfer function has been determined to recover from the pulse integral the number of photo-electrons.
@@ 4.2.2 The optical point spread function of the telescope
The PSF represents the optical quality of the telescope. We define the PSF as the angular diameter of a circle containing 80% of the photons emitted by a point source at infinity, denoted by D80. The centre of the circle is chosen to be the centre of gravity of the photon distribution. Given that the spot profile at the focal plane is not necessarily symmetrical, we define D80 by starting from the centre of gravity of the photon distribution and integrating the signal in larger and larger circles until 80% of the total emission is reached. For the camera field of view of 9, the PSF needs to be determined for light rays coming from a source located at $\sqrt{0.8}\times 4.5^{\circ}$, i.e. at 4 off-axis. The PSF simulation for the various field of view of the telescope using two independent simulation tools is shown in Fig. 4. The simulation is compared with the real measured PSF of the telescope in Sec. 2 and it is shown in 2 dimensions in Fig. 30.
@@ 4.3.1 Prototype structure tests
To fully identify the spectrum of natural modes of the telescope structure, additional modal experiments were conducted using piezoelectric accelerometers by a specialised team from AGH in Krakow. Two accelerometers were used – one at the bottom of the dish support structure and another on the back of the camera box. The structure was excited with a PCB 086D20 impact hammer at various points on the frame and different directions. Points of excitation were approximately the same across different elevation angles ($-10^{\circ}$ – close to parking position, but telescope not locked in docking station, $-13^{\circ}$ – telescope locked, $45^{\circ}$ and $90^{\circ}$), except near the camera box at $45^{\circ}$ and $90^{\circ}$ angles. Nineteen modal experiments were performed, varying the excitation points (dish support, mast, camera box, counterweight) and directions ($x$, $y$, $z$). Some measurements were conducted with the drives activated (typical during observation) and others with the drives turned off and brakes engaged (as in the parking position). The reference frame was centred at the base of the telescope tower, with the $x$-axis aligned along the telescope mast and the $y$- and $z$-axes oriented vertically. The analysis was based on spectral transfer functions and coherence functions. The spectral transfer function provides information about the natural frequencies excited during an experiment, while the coherence function indicates the correlation between the excitation and the response, ranging from 0 to 1. Results are considered valid when the coherence function values exceed 0.8. Frequencies analysed ranged from 1.9 Hz to 32 Hz.
@@ 4.4.2 Telescope optical efficiency
We calculate the fraction of the Cherenkov light signal that can be detected by the telescopes using the calculated total efficiency from 290 nm to 700 nm: 15.1% and 16.5% for the prototype and the second telescope. The same efficiencies for the NSB are 12.2% and 13.6%. Two expressions of SNRs are used, the first one only considers the NSB as source of noise while the second one account for fluctuations of the signal itself as part of the noise. The resulting SNRs are shown in Fig. 35 (bottom). For comparison, the SNRs labelled "No cut-off" displays the value computed if the entrance window was identical to the one of the Large-Sized Telescope of CTA. This window is made out of a special PMMA, called Shinkolite, with enhanced transmissivity in the UV region. One can clearly see the positive impact of the low-pass optical filter, which stabilises the SNR ratio above 540 nm. This is observed independently of the SNR expression.
Figure 35: Top: Efficiency of the optical elements dedicated to transmission of light from the mirror dish. The total efficiencies of the 2 SST-1Ms are 15.1% and 16.5%. Bottom: SNRs for the two telescopes as well the one computed if the window would be identical to the one mounted on the Large-Sized Telescope that does not have any wavelength cut-off.
@@ 4.4.3 Sensor characteristics
Gain [ADC] Dark count rate [MHz] Optical cross talk [%] $\sigma_{pe}$ [ADC] $\sigma_{el}$ [ADC]
@@
Table 9: Summary of mean and standard deviation of the distributions of the main SiPM parameters
@@ 4.6 Operation at the Ondřejov site
Figure 43: The location of the two SST-1Ms at the Ondřejov observatory (left). The layout of the telescope surroundings. The telescope is secured with a fence accessible via a gate. The camera servers and all necessary control electronics are located in a wooden booth at the top right of the image on the right image. The DAQ and slow control are located in the control cabinet next to the chiller. Imagery ©Airbus, CNES/ Airbus, Maxar Technologies, Maps data ©2025
@@ 5 Summary
The high-performance, large-area SiPMs with dedicated slow control, their lightweight (8.6 t) and compact mechanical telescope structure and its robotic operation require minimal maintenance in harsh environment, and make the SST-1M suitable for deployment at high-altitude sites above 4000 m a.s.l. The SiPM market and their applications have widely expanded in the automotive, biology, medical imaging and LIDAR fields. Major companies such as Broadcom, Hamamatsu, Ketek, On-Semi, Advansid, and FBK are highly competitive and rapidly reducing costs.
@@ Acknowledgments
This publication was created as part of the projects funded in Poland by the Minister of Science based on agreements number 2024/WK/03 and DIR/WK/2017/2022/12-3.
The construction, calibration, software control and support for operation of the SST-1M cameras is supported by SNF (grants CRSII2_141877, 20FL21_154221, CRSII2_160830, _166913), by the Boninchi Foundation and by the Université de Genève, Faculté de Sciences, Départment de Physique Nucléaire. The Czech partner institutions acknowledge support of the infrastructure and research projects by Ministry of Education, Youth and Sports of the Czech Republic and regional funds of the European Union, MEYS LM2023047 and EU/MEYS CZ.02.01.01/00/22 008/0004632, and Czech Science Foundation, GACR 23-05827S.
Appendix A Telescope and infrastructure diagrams